OpenStudy (anonymous):

Determine the slope between the points (-1, 0) and (-2, 3) A. 3 B. -1/3 C. -3 D. 1/3

OpenStudy (anonymous):

C

OpenStudy (anonymous):

(0-3)/(-1-(-2))=(-3)/1)=-3 (y2-y1)/(x2-x1)